# Algorithms for Generating Small Random Samples

### v2

@@ 4.1 Sampling Pairs of Distinct Integers
This first set of results in Table 3 explores the difference in runtime performance of the various approaches to sampling random pairs of distinct integers from the half open interval $[0,n)$. It is clear from the results that the RandomPair (Algorithm 1) of this paper dominates all of the others. Its average runtime is a low constant, and does not vary as $n$ increases, at least not to any meaningful degree. If it appears that average time is possibly decreasing as $n$ increases for this algorithm, it may be, although it is clearly negligible. This is because as $n$ increases, the probability of the assignment statement on line 4 of Algorithm 1 decreases. The insertion sampling [2] average runtime is likewise approximately constant, though a higher constant. Note that insertion sampling’s theoretical runtime is $O(k^{2})$, which is essentially constant for any fixed $k$. Pool sampling [5] and reservoir sampling [11] both have an average runtime that increases linearly in $n$. Reservoir sampling’s average runtime is much higher than pool sampling, however, because reservoir sampling also generates $(n-k)$ random integers to sample $k$ integers from a set of $n$ integers, while pool sampling only needs $k$ random integers to accomplish this. Thus, reservoir is an especially poor choice for low $k$ relative to $n$.
@@
$n$ RandomPair (Algorithm 1) insertion [2] pool [5] reservoir [11]
$n=16$ $15.899\pm 0.017$ $24.125\pm 0.271$ $55.035\pm 0.096$ $178.878\pm 4.264$
$n=64$ $15.580\pm 0.128$ $23.863\pm 0.038$ $89.234\pm 2.463$ $633.099\pm 2.630$
$n=256$ $15.485\pm 0.044$ $24.175\pm 0.106$ $295.983\pm 3.101$ $2087.132\pm 25.733$
$n=1024$ $15.451\pm 0.023$ $24.153\pm 0.088$ $1199.580\pm 8.026$ $7931.636\pm 25.100$
@@ 4.2 Sampling Triples of Distinct Integers
We now consider the results in Table 5 for sampling random triples of distinct integers. As in the case of sampling pairs of distinct integers, we find that the algorithm of this paper, RandomTriple (Algorithm 2) requires a constant time on average. The insertion sampling algorithm likewise has an average time that is constant, although a higher constant than Algorithm 2. Both pool sampling and reservoir sampling require time that increases linearly in $n$, with reservoir sampling’s times significantly higher than pool sampling due to generating significantly more random numbers during the sampling process for low $k$.
@@
$n$ RandomTriple (Algorithm 2) insertion [2] pool [5] reservoir [11]
$n=16$ $24.795\pm 0.201$ $41.775\pm 0.302$ $59.406\pm 0.707$ $179.879\pm 1.738$
$n=64$ $24.224\pm 0.054$ $41.881\pm 0.167$ $90.707\pm 0.762$ $653.295\pm 1.533$
$n=256$ $23.100\pm 0.032$ $45.836\pm 0.325$ $289.158\pm 0.614$ $2115.373\pm 141.514$
$n=1024$ $22.889\pm 0.098$ $41.244\pm 0.068$ $1174.421\pm 12.878$ $8575.434\pm 6.903$
In Table 6 we further explore the effects of allocating a new array for each sampled triple, versus using a preallocated array, versus returning a Java record. The results in the triple of integers case does not follow the same pattern as that of the pairs of integers case. For triples of distinct integers, the preallocated array case is the fastest, whereas for pairs, the Java record case was the fastest. For triples, the difference in performance between allocating a new array versus a Java record is less clear.
@@ 5 Conclusions
In this report, we presented constant runtime algorithms for sampling pairs and triples of distinct integers from the interval $[0,n)$. We conducted benchmarking experiments with our Java implementations using JMH that demonstrate that these sampling algorithms are substantially faster than using existing general purpose algorithms for sampling $k$ integers from $[0,n)$. For example, reservoir sampling [11] and pool sampling [5] both require linear time, with reservoir sampling being especially computationally expensive for low $k$ since it also requires $n-k$ random integers for sample size $k$. Although insertion sampling [2] has a constant runtime for any fixed $k$ (e.g., runtime increases with $k$ but not with $n$), the special purpose algorithms for the two specific cases of $k=2$ and $k=3$ presented in this paper are significantly faster.
Our implementations of all of the algorithms of this paper (both new and existing algorithms) are available in the open source Java library $\rho\mu$. The experiments of this paper are also available on GitHub to assist with reproducibility.

### v3

@@ 4.1 Sampling Pairs of Distinct Integers
This first set of results in Table 3 explores the difference in runtime performance of the various approaches to sampling random pairs of distinct integers from the half open interval $[0,n)$. It is clear from the results that the RandomPair (Algorithm 1) of this paper dominates all of the others. Its average runtime is a low constant, and does not vary as $n$ increases, at least not to any significant degree. If it appears that average time is possibly decreasing as $n$ increases for this algorithm, it may be, although it is clearly negligible. This is because as $n$ increases, the probability of the conditional assignment on line 3 of Algorithm 1 decreases. The insertion sampling 4 average runtime is likewise approximately constant, though a higher constant. Note that insertion sampling’s theoretical runtime is $O(k^{2})$, which is essentially constant for any fixed $k$. Pool sampling 3 and reservoir (R) sampling 1 both have an average runtime that increases linearly in $n$. Reservoir (R) sampling’s average runtime is much higher than pool sampling, however, because reservoir (R) sampling also generates $(n-k)$ random integers to sample $k$ integers from a set of $n$ integers, while pool sampling only needs $k$ random integers to accomplish this. Thus, reservoir (R) is an especially poor choice for low $k$ relative to $n$. Reservoir (L) sampling 2 is the slowest for low $n$, but it overtakes reservoir (R) as $n$ increases, and should surpass the performance of pool sampling if we further increase $n$. Its runtime increases with $n$, but at a slower rate $O(k(1+\ln(n/k)))$ than reservoir (R) and pool sampling.
@@
$n$ RandomPair (Algorithm 1) insertion 4 pool 3 reservoir (R) 1 reservoir (L) 2
$n=16$ $16.0\pm 0.5$ $24.1\pm{<}0.1$ $54.9\pm 0.3$ $178.2\pm 1.0$ $465.8\pm 14.5$
$n=64$ $15.6\pm 0.1$ $24.1\pm 0.1$ $89.3\pm 1.9$ $633.4\pm 9.7$ $721.3\pm 19.5$
$n=256$ $15.7\pm 0.6$ $24.3\pm{<}0.1$ $298.8\pm 0.8$ $2062.5\pm 9.7$ $981.3\pm 8.4$
$n=1024$ $15.7\pm{<}0.1$ $23.6\pm 0.1$ $1209.9\pm 16.2$ $7918.2\pm 12.9$ $1242.7\pm 1.4$
@@ 4.2 Sampling Triples of Distinct Integers
We now consider the results in Table 5 for sampling random triples of distinct integers. As in the case of sampling pairs of distinct integers, we find that the algorithm of this paper, RandomTriple (Algorithm 2) requires a constant time on average. The insertion sampling algorithm likewise has an average time that is constant, although a higher constant than Algorithm 2. Both pool sampling and reservoir (R) sampling require time that increases linearly in $n$, with reservoir sampling’s times significantly higher than pool sampling due to generating significantly more random numbers during the sampling process for low $k$. Reservoir (L) sampling is also significantly slower than Algorithm 2.
@@
$n$ RandomTriple (Algorithm 2) insertion 4 pool 3 reservoir (R) 1 reservoir (L) 2
$n=16$ $23.6\pm{<}0.1$ $41.8\pm{<}0.1$ $59.4\pm 0.2$ $179.3\pm 0.3$ $583.4\pm 0.8$
$n=64$ $22.4\pm 0.3$ $41.1\pm 0.2$ $91.3\pm 0.3$ $652.0\pm 0.7$ $1026.6\pm 3.4$
$n=256$ $22.4\pm{<}0.1$ $45.6\pm 0.1$ $286.0\pm 1.1$ $2113.4\pm 23.6$ $1459.3\pm 103.8$
$n=1024$ $21.8\pm 0.1$ $42.3\pm 0.3$ $1177.5\pm 3.7$ $8564.1\pm 64.3$ $1867.6\pm 292.3$
In Table 6 we further explore the effects of allocating a new array for each sampled triple, versus using a preallocated array, versus returning a Java record. The results in the triple of integers case follow the same pattern as that of the pairs of integers case. For triples of distinct integers, returning a Java record is about 10% faster than using a preallocated array for the result, which in turn is about 10% faster than returning a new array.
@@ 5 Conclusions
In this article, we presented constant runtime algorithms for sampling pairs and triples of distinct integers from the interval $[0,n)$. We conducted benchmarking experiments with our Java implementations using JMH that demonstrate that these sampling algorithms are substantially faster than using existing general purpose algorithms for sampling $k$ integers from $[0,n)$. For example, reservoir (R) sampling 1 and pool sampling 3 both require linear time, with reservoir (R) sampling being especially computationally expensive for low $k$ since it also requires $n-k$ random integers for sample size $k$. Reservoir (L) sampling 2 is superior to reservoir (R) for larger $n$, but it is still very slow relative to the constant time algorithms for the special cases of $k=2$ and $k=3$. Although insertion sampling 4 has a constant runtime for any fixed $k$ (e.g., runtime increases with $k$ but not with $n$), the special purpose algorithms for the two specific cases of $k=2$ and $k=3$ presented in this paper are significantly faster.
The structure of Algorithms 1 and 2 for sampling pairs and triples of distinct integers suggest a pattern for forming a special purpose algorithm for any fixed $k$, such as the example provided for $k=4$ in Algorithm 3. Although it is unlikely practical or efficient for larger $k$ in software, it is possible that a circuit implementation of such a sampling network may have similar benefits as sorting networks.
Our implementations of the algorithms for small random samples and three of the general purpose sampling algorithms are available in the open source Java library $\rho\mu$. The experiments of this paper, including the fourth general purpose sampling algorithm, are also available on GitHub to assist with reproducibility.
[…]
*Conflict of interest
The author declares no potential conflict of interests.